OpenStudy (anonymous):

solve 3x+5y=22 4x+3y=11 show work

OpenStudy (anonymous):

x first equn. by 4 and second by 3 12x + 20y = 88 12x +9y = 33 11y = 55 y =5 first equn. 3x + 5*5 = 22 3x = -3 x = -1 y=5,x=-1

OpenStudy (anonymous):

x=-1 y=5